# Leganés, Spain 22affiliationtext: Instituto de Investigación Sanitaria Gregorio Marañón Madrid, Spain

### v2

@@ 2 Behavioral Dataset
    $1$ represents weekend data, whereas 00 represents weekday data.
@@ 3.1 I-TSFM: A Transformer Approach
As a result, the temporal sequence used as input to the model has dimensions $L\times S$, where $L$ represents the total sequence length and $S$, the probability for each of the hidden states. The input for the transformer encoder and decoder were reshaped into observation batches with dimensions $\left[B,\,L_{e},\,S\right]$ and $\left[B,\,L_{d},\,S\right]$, where $B$ denotes the batch size, $L_{e}$ and $L_{d}$ the encoder and decoder sequence lengths, respectively, and $S$ is the feature dimension. The architecture of the transformer model is described in Appendix B of the Supplementary Material.
@@ 3.2 Our VQ-TSFM: A Quantized VAE Approach
Let $\mathbf{x}_{d}^{(i)}\in\mathbb{R}^{T}$ represent the time-series data vector of length $T$ for patient $i$ and variable $d$, where each component corresponds to a data entry in a sampled time instant and $d\in\{1,\ldots,D\}$. Recall that the set of possible variables is summarized in Table 1. Let $\mathbf{m}_{d}^{(i)}\in\{0,1\}^{T}$ denote a binary mask vector where each entry indicates whether the corresponding entry is observed (entry value equal to $1$) or missing (entry value equal to 00). The corrupted signal, after applying the binary mask, $\mathbf{m}_{d}^{(i)}$, is defined as:
@@ Appendix A Data Preprocessing for the VQ-TSFM
Several variables were excluded from the analysis due to inconsistencies in missing data reporting. For instance, features such as the variables measuring the minimum/maximum/average heart rate used a placeholder value of $-1$ to indicate missing data, whereas other variables adhered to the standard Numpy convention of using NaN. Date-related variables also required normalization to a consistent format. Additionally, certain variables contained erroneous or outlier values, likely due to faulty sensors or other external factors, as discussed in Section 2. While it was not possible to completely eliminate all erroneous entries due to the absence of key contextual variables, we removed the majority of manifestly inaccurate data points. For example, the Sleep Duration variable is known to be device-dependent, with different vendors applying varying algorithms to detect sleep patterns. Similarly, the Total Steps variable can be influenced by non-step movements, such as hand gestures, whereas the App Usage variable is constrained by vendor-specific limitations. The Location Clusters variable, being derived from external algorithms that process raw geolocation data, also exhibited potential inaccuracies.
@@ Appendix B Informer Architectural Details
The Informer approach described in Section 3.1 follows an encoder-decoder transformer architecture. Appendix B shows the details of each layer of such model. The input for the encoder and decoder were reshaped into batches of dimensions [$B$, $L_{e}$, $S$] and [$B$, $L_{d}$, $S$], respectively, where $B$ denotes the batch size, $L_{e}$ and $L_{d}$ the encoder and decoder sequence lengths, and $S$ is the feature dimension.
@@
Input (Signal) $[B,L_{e},S]$ Encoder input (sequence of features)
Embedding Transformation $[B,L_{e},S]$ $[B,L_{e},d_{\text{embedding}}]$ $\mathbf{X}_{\text{enc}}\in\mathbb{R}^{L_{e}\times S}\to\mathbb{R}^{L_{e}\times d_{\text{embedding}}}$
Positional Encoding Addition $[B,L_{e},d_{\text{embedding}}]$ $[B,L_{e},d_{\text{embedding}}]$ Add sinusoidal positional encoding
[…]
Add and Normalization $[B,L_{e},d_{\text{embedding}}]$ $[B,L_{e},d_{\text{embedding}}]$ Residual add + LayerNorm
Feed-Forward Layer $[B,L_{e},d_{\text{embedding}}]$ $[B,L_{e},d_{\text{embedding}}]$ FC + ReLU
Add and Normalization $[B,L_{e},d_{\text{embedding}}]$ $[B,L_{e},d_{\text{embedding}}]$ Residual add + LayerNorm
Conv1D + ELU Activation $[B,L_{e},d_{\text{embedding}}]$ $[B,L_{e},d_{\text{embedding}}]$ 1D conv (kernel=3, stride=1) + ELU
@@
Input (Signal) $[B,L_{d},S]$ Decoder input (sequence of features)
Embedding Transformation $[B,L_{d},S]$ $[B,L_{d},d_{\text{embedding}}]$ $\mathbf{X}_{\text{dec}}\in\mathbb{R}^{L_{d}\times S}\to\mathbb{R}^{L_{d}\times d_{\text{embedding}}}$
Positional Encoding Addition $[B,L_{d},d_{\text{embedding}}]$ $[B,L_{d},d_{\text{embedding}}]$ Add sinusoidal positional encoding
[…]
Add and Normalization $[B,L_{d},d_{\text{embedding}}]$ $[B,L_{d},d_{\text{embedding}}]$ Residual add + LayerNorm
Cross-Attention $[B,L_{d},d_{\text{embedding}}]$ $[B,L_{d},d_{\text{embedding}}]$ Attends to encoder outputs
Add and Normalization $[B,L_{d},d_{\text{embedding}}]$ $[B,L_{d},d_{\text{embedding}}]$ Residual add + LayerNorm
Feed-Forward Layer $[B,L_{d},d_{\text{embedding}}]$ $[B,L_{d},d_{\text{embedding}}]$ FC + ReLU
Add and Normalization $[B,L_{d},d_{\text{embedding}}]$ $[B,L_{d},d_{\text{embedding}}]$ Residual add + LayerNorm
Linear Projection $[B,L_{d},d_{\text{embedding}}]$ $[B,L_{d},S]$ Project to output space
Softmax $[B,L_{d},S]$ $[B,L_{d},S]$ Probability distribution
@@ Appendix C VQ-TSFM Architectural Details
The architectures for the base model VQ-TSFM and its extensions E1 and E2 are illustrated in Figures 7(a), 7(b), and 7(c), respectively. Throughout the network, spatial length was preserved to ensure that each time step—representing daily patient states—was captured in the embeddings.
@@
Input (Signal) $[B,\,F,\,L]$ Model input (signal)
Conv-Block 1 $[B,\,F,\,L]$ $[B,\,F,\,L]$ Convolutional blocka
[…]
@@
[…]
@@
Input (Signal) $[B,\,F,\,L]$ –– Model input (signal)
Input (Mask) $[B,\,M,\,L]$ Model input (mask)
Mask-Conv-Block 1 $[B,\,M,\,L]$ $[B,\,M,\,L]$ Convolutional blocka
Mask-Conv-Block 2 $[B,\,M,\,L]$ $[B,\,M,\,L]$
Concatenation (Signal + Mask) $[B,\,F,\,L]$, $[B,\,M,\,L]$ $[B,\,F+M,\,L]$ Note: $F=M$
Conv-Block 1 $[B,\,F+M,\,L]$ $[B,\,F,\,L]$
[…]
@@
[…]
@@
Input (Signal) $[B,\,F,\,L]$ –– Model input (signal)
Input (Mask) $[B,\,M,\,L]$ Model input (mask)
Mask-Conv-Block 1 $[B,\,M,\,L]$ $[B,\,M,\,L]$ Convolutional blocka
Mask-Conv-Block 2 $[B,\,M,\,L]$ $[B,\,M,\,L]$
Concatenation (Signal + Mask) $[B,\,F,\,L]$, $[B,\,M,\,L]$ $[B,\,F+M,\,L]$ Note: $F=M$
Conv-Block 1 $[B,\,F+M,\,L]$ $[B,\,F,\,L]$
[…]
@@
[…]
Input (Mask) $[B,\,M,\,L]$ –– Model input (mask)
Mask-Conv-Block 1 $[B,\,M,\,L]$ $[B,\,M,\,L]$
Mask-Conv-Block 2 $[B,\,M,\,L]$ $[B,\,M,\,L]$ —–
[…]
Concatenation (Quantized Signal + Mask) $[B,\,F,\,L]$, $[B,\,M,\,L]$ $[B,\,F+M,\,L]$ Note: $F=M$
Fine-tuning-Block 1 $[B,\,F+M,\,L]$ $[B,\,F+M,\,L]$
Fine-tuning-Block 2 $[B,\,F+M,\,L]$ $[B,\,F,\,L]$ —–
Fine-tuning-Block 3 $[B,\,F,\,L]$ $[B,\,F,\,L]$ —–
Fine-tuning-Block 4 $[B,\,F,\,L]$ $[B,\,F,\,L]$ Last deconvolutional blockb
@@ Appendix E CPD Algorithm Details and Ablation Study
  • Number of profiles, $K$. While not a hyperparameter of the CPD stage (but rather involved in the previous profiling step), the number of possible profiles is a crucial setting in the downstream task. Too few profiles will fail to capture the distinct behavior patterns, but too many may introduce noisy profiles modeled with low confidence that impede the correct performance of the CPD. The value of $K$ in the heterogeneous mixture model was analyzed (8(a)) and chosen to be 10, the one yielding the best results. Notice that the VQ-TSFM model used to compare may use a different number of profiles. In fact, Figure 10 suggests that $K=20$ is the most suitable value in the VQ-TSFM context. On the other hand, the Informer approach does not have this parameter because its latent representation are real-valued embeddings.
  • Number of samples in multinomial distribution, $S$. In the multinomial approach, $S$ represents the number of samples that are drawn from the posterior distribution of the latent variables at each time step. A larger value will adapt better to the latent profiles but also complicates the detection task of the CPD. As evidenced by the results shown in 8(b), $S=5$ provides the most suitable sample size for our task.
@@
  • Size of the temporal window, $W$. The CPD model focuses on a temporal frame to assess whether its predictions are successful or not. For example, for each true event, a true positive is returned if an alarm was given by the model within the temporal window previous to that event. If the CPD did not predict any change, then a false negative is counted. This window parameter allows therefore to select how long in advance we aim to predict suicide events. We chose a prediction period of one week ($W=7$ days), which obtained a high AUC in our analysis (see 8(c)) and is brief enough to serve as short-term prediction.
@@
    • MAP ratio (default) → based on the MAP estimates of the run length, an alarm is returned if the ratio of current $r_{t}$ over the previous day $r_{t-1}$ is below the threshold: rtrt-1 < τ
    • MAP difference → based on the MAP estimates of the run length, an alarm is returned if the difference between current $r_{t}$ and previous $r_{t-1}$ is above the threshold: r_t - r_t-1 > τ
    • Cumulative sum → based on the cumulative probability of the run length of previous days (within the specified window of size $W$), an alarm is returned if this sum is above the threshold: ∑_i=0^W r_t-i > τ

### v3

# A Vector-Quantized Foundation Model for Patient Behavior Monitoring

@@ 2 Behavioral Dataset
    11 represents weekend data, whereas 0 represents weekday data.
@@ 3.1 I-TSFM: A Transformer Approach
As a result, the temporal sequence used as input to the model has dimensions $L\times S$, where $L$ represents the total sequence length and $S$, the probability for each of the hidden states. The input for the transformer encoder and decoder were reshaped into observation batches with dimensions $\left[B,\,L_{e},\,S\right]$ and $\left[B,\,L_{d},\,S\right]$, where $B$ denotes the batch size, $L_{e}$ and $L_{d}$ the encoder and decoder sequence lengths, respectively, and $S$ is the feature dimension. The architecture of the transformer model is described in Table 5 of the Supplementary Material.
@@ 3.2 Our VQ-TSFM: A Quantized VAE Approach
Let $\mathbf{x}_{d}^{(i)}\in\mathbb{R}^{T}$ represent the time-series data vector of length $T$ for patient $i$ and variable $d$, where each component corresponds to a data entry in a sampled time instant and $d\in\{1,\ldots,D\}$. Recall that the set of possible variables is summarized in Table 1. Let $\mathbf{m}_{d}^{(i)}\in\{0,1\}^{T}$ denote a binary mask vector where each entry indicates whether the corresponding entry is observed (entry value equal to $1$) or missing (entry value equal to 0). The corrupted signal, after applying the binary mask, $\mathbf{m}_{d}^{(i)}$, is defined as:
@@ Appendix A Data Preprocessing for the VQ-TSFM
Several variables were excluded from the analysis due to inconsistencies in missing data reporting. For instance, features such as the variables measuring the minimum/maximum/average heart rate used a placeholder value of 1-1 to indicate missing data, whereas other variables adhered to the standard Numpy convention of using NaN. Date-related variables also required normalization to a consistent format. Additionally, certain variables contained erroneous or outlier values, likely due to faulty sensors or other external factors, as discussed in Section 2. While it was not possible to completely eliminate all erroneous entries due to the absence of key contextual variables, we removed the majority of manifestly inaccurate data points. For example, the Sleep Duration variable is known to be device-dependent, with different vendors applying varying algorithms to detect sleep patterns. Similarly, the Total Steps variable can be influenced by non-step movements, such as hand gestures, whereas the App Usage variable is constrained by vendor-specific limitations. The Location Clusters variable, being derived from external algorithms that process raw geolocation data, also exhibited potential inaccuracies.
@@ Appendix B Informer Architectural Details
The Informer approach described in Section 3.1 follows an encoder-decoder transformer architecture. Table 5 shows the details of each layer of such model. The input for the encoder and decoder were reshaped into batches of dimensions [$B$, $L_{e}$, $S$] and [$B$, $L_{d}$, $S$], respectively, where $B$ denotes the batch size, $L_{e}$ and $L_{d}$ the encoder and decoder sequence lengths, and $S$ is the feature dimension.
@@
Input (Signal) $B,L_{e},S]$ Encoder input (sequence of features)
Embedding Transformation $B,L_{e},S] [$$B,L_{e},d_{\text{embedding}}]$ $\mathbf{X}_{\text{enc}}\in\mathbb{R}^{L_{e}\times S}\to\mathbb{R}^{L_{e}\times d_{\text{embedding}}}$
Positional Encoding Addition $B,L_{e},d_{\text{embedding}}] [$$B,L_{e},d_{\text{embedding}}]$ Add sinusoidal positional encoding
3 $\times$ Attention Blocks
ProbSparse Self-Attention $B,L_{e},d_{\text{embedding}}] [$$B,L_{e},d_{\text{embedding}}]$ Efficient self-attention (O($L\log L$))
Add and Normalization $B,L_{e},d_{\text{embedding}}] [$$B,L_{e},d_{\text{embedding}}]$ Residual add + LayerNorm
Feed-Forward Layer $B,L_{e},d_{\text{embedding}}] [$$B,L_{e},d_{\text{embedding}}]$ FC + ReLU
Add and Normalization $B,L_{e},d_{\text{embedding}}] [$$B,L_{e},d_{\text{embedding}}]$ Residual add + LayerNorm
Conv1D + ELU Activation $B,L_{e},d_{\text{embedding}}] [$$B,L_{e},d_{\text{embedding}}]$ 1D conv (kernel=3, stride=1) + ELU
@@ 3 $\times$ Attention Blocks
Input (Signal) $B,L_{d},S]$ Decoder input (sequence of features)
Embedding Transformation $B,L_{d},S] [$$B,L_{d},d_{\text{embedding}}]$ $\mathbf{X}_{\text{dec}}\in\mathbb{R}^{L_{d}\times S}\to\mathbb{R}^{L_{d}\times d_{\text{embedding}}}$
Positional Encoding Addition $B,L_{d},d_{\text{embedding}}] [$$B,L_{d},d_{\text{embedding}}]$ Add sinusoidal positional encoding
3 $\times$ Attention Blocks
Masked Self-Attention $B,L_{d},d_{\text{embedding}}] [$$B,L_{d},d_{\text{embedding}}]$ Prevents attending to future tokens
Add and Normalization $B,L_{d},d_{\text{embedding}}] [$$B,L_{d},d_{\text{embedding}}]$ Residual add + LayerNorm
Cross-Attention $B,L_{d},d_{\text{embedding}}] [$$B,L_{d},d_{\text{embedding}}]$ Attends to encoder outputs
Add and Normalization $B,L_{d},d_{\text{embedding}}] [$$B,L_{d},d_{\text{embedding}}]$ Residual add + LayerNorm
Feed-Forward Layer $B,L_{d},d_{\text{embedding}}] [$$B,L_{d},d_{\text{embedding}}]$ FC + ReLU
Add and Normalization $B,L_{d},d_{\text{embedding}}] [$$B,L_{d},d_{\text{embedding}}]$ Residual add + LayerNorm
Linear Projection $B,L_{d},d_{\text{embedding}}] [$$B,L_{d},S]$ Project to output space
Softmax $B,L_{d},S] [$$B,L_{d},S]$ Probability distribution
@@ Appendix C VQ-TSFM Architectural Details
The architectures for the base model VQ-TSFM and its extensions E1 and E2 are illustrated in Figures 8(a), 8(b), and 8(c), respectively. Throughout the network, spatial length was preserved to ensure that each time step—representing daily patient states—was captured in the embeddings.
@@
Input (Signal) [B,F,L][B,\,F,\,L] Model input (signal)
Conv-Block 1 [B,F,L][B,\,F,\,L] [B,F,L][B,\,F,\,L] Convolutional blocka
Conv-Block 2 [B,F,L][B,\,F,\,L] [B, 2F,L][B,\,2F,\,L]
Conv-Block 3 [B, 2F,L][B,\,2F,\,L] [B, 4F,L][B,\,4F,\,L]
Conv-Block 4 [B, 4F,L][B,\,4F,\,L] [B, 8F,L][B,\,8F,\,L]
@@
Deconv-Block 1 [B, 8F,L][B,\,8F,\,L] [B, 6F,L][B,\,6F,\,L]
Deconv-Block 2 [B, 6F,L][B,\,6F,\,L] [B, 4F,L][B,\,4F,\,L] —–
Deconv-Block 3 [B, 4F,L][B,\,4F,\,L] [B, 4F,L][B,\,4F,\,L] —–
Deconv-Block 4 [B, 4F,L][B,\,4F,\,L] [B, 2F,L][B,\,2F,\,L] —–
Deconv-Block 5 [B, 2F,L][B,\,2F,\,L] [B,F,L][B,\,F,\,L] Last deconvolutional blockb
@@
Input (Signal) [B,F,L][B,\,F,\,L] –– Model input (signal)
Input (Mask) [B,M,L][B,\,M,\,L] Model input (mask)
Mask-Conv-Block 1 [B,M,L][B,\,M,\,L] [B,M,L][B,\,M,\,L] Convolutional blocka
Mask-Conv-Block 2 [B,M,L][B,\,M,\,L] [B,M,L][B,\,M,\,L]
Concatenation (Signal + Mask) [B,F,L][B,\,F,\,L], [B,M,L][B,\,M,\,L] [B,F+M,L][B,\,F+M,\,L] Note: $F=M$
Conv-Block 1 [B,F+M,L][B,\,F+M,\,L] [B,F,L][B,\,F,\,L]
Conv-Block 2 [B,F,L][B,\,F,\,L] [B, 2F,L][B,\,2F,\,L]
Conv-Block 3 [B, 2F,L][B,\,2F,\,L] [B, 4F,L][B,\,4F,\,L]
Conv-Block 4 [B, 4F,L][B,\,4F,\,L] [B, 4F,L][B,\,4F,\,L]
Conv-Block 5 [B, 4F,L][B,\,4F,\,L] [B, 6F,L][B,\,6F,\,L]
Conv-Block 6 [B, 6F,L][B,\,6F,\,L] [B, 8F,L][B,\,8F,\,L]
@@
Deconv-Block 1 [B, 8F,L][B,\,8F,\,L] [B, 6F,L][B,\,6F,\,L]
Deconv-Block 2 [B, 6F,L][B,\,6F,\,L] [B, 4F,L][B,\,4F,\,L] —–
Deconv-Block 3 [B, 4F,L][B,\,4F,\,L] [B, 4F,L][B,\,4F,\,L] —–
Deconv-Block 4 [B, 4F,L][B,\,4F,\,L] [B, 2F,L][B,\,2F,\,L] —–
Deconv-Block 5 [B, 2F,L][B,\,2F,\,L] [B,F,L][B,\,F,\,L] Last deconvolutional blockb
@@
Input (Signal) [B,F,L][B,\,F,\,L] –– Model input (signal)
Input (Mask) [B,M,L][B,\,M,\,L] Model input (mask)
Mask-Conv-Block 1 [B,M,L][B,\,M,\,L] [B,M,L][B,\,M,\,L] Convolutional blocka
Mask-Conv-Block 2 [B,M,L][B,\,M,\,L] [B,M,L][B,\,M,\,L]
Concatenation (Signal + Mask) [B,F,L][B,\,F,\,L], [B,M,L][B,\,M,\,L] [B,F+M,L][B,\,F+M,\,L] Note: $F=M$
Conv-Block 1 [B,F+M,L][B,\,F+M,\,L] [B,F,L][B,\,F,\,L]
Conv-Block 2 [B,F,L][B,\,F,\,L] [B, 2F,L][B,\,2F,\,L]
Conv-Block 3 [B, 2F,L][B,\,2F,\,L] [B, 4F,L][B,\,4F,\,L]
Conv-Block 4 [B, 4F,L][B,\,4F,\,L] [B, 4F,L][B,\,4F,\,L]
Conv-Block 5 [B, 4F,L][B,\,4F,\,L] [B, 6F,L][B,\,6F,\,L]
Conv-Block 6 [B, 6F,L][B,\,6F,\,L] [B, 8F,L][B,\,8F,\,L]
@@
Input (Quantized Signal) [B, 8F,L][B,\,8F,\,L] –– Model input (quantized signal)
Input (Mask) [B,M,L][B,\,M,\,L] –– Model input (mask)
Mask-Conv-Block 1 [B,M,L][B,\,M,\,L] [B,M,L][B,\,M,\,L]
Mask-Conv-Block 2 [B,M,L][B,\,M,\,L] [B,M,L][B,\,M,\,L] —–
Deconv-Block 1 [B, 8F,L][B,\,8F,\,L] [B, 6F,L][B,\,6F,\,L]
Deconv-Block 2 [B, 6F,L][B,\,6F,\,L] [B, 4F,L][B,\,4F,\,L] —–
Deconv-Block 3 [B, 4F,L][B,\,4F,\,L] [B, 4F,L][B,\,4F,\,L] —–
Deconv-Block 4 [B, 4F,L][B,\,4F,\,L] [B, 2F,L][B,\,2F,\,L] —–
Deconv-Block 5 [B, 2F,L][B,\,2F,\,L] [B,F,L][B,\,F,\,L] —–
Concatenation (Quantized Signal + Mask) [B,F,L][B,\,F,\,L], [B,M,L][B,\,M,\,L] [B,F+M,L][B,\,F+M,\,L] Note: $F=M$
Fine-tuning-Block 1 [B,F+M,L][B,\,F+M,\,L] [B,F+M,L][B,\,F+M,\,L]
Fine-tuning-Block 2 [B,F+M,L][B,\,F+M,\,L] [B,F,L][B,\,F,\,L] —–
Fine-tuning-Block 3 [B,F,L][B,\,F,\,L] [B,F,L][B,\,F,\,L] —–
Fine-tuning-Block 4 [B,F,L][B,\,F,\,L] [B,F,L][B,\,F,\,L] Last deconvolutional blockb
@@ Appendix E CPD Algorithm Details and Ablation Study
  • Number of profiles, $K$. While not a hyperparameter of the CPD stage (but rather involved in the previous profiling step), the number of possible profiles is a crucial setting in the downstream task. Too few profiles will fail to capture the distinct behavior patterns, but too many may introduce noisy profiles modeled with low confidence that impede the correct performance of the CPD. The value of $K$ in the heterogeneous mixture model was analyzed (9(a)) and chosen to be 10, the one yielding the best results. Notice that the VQ-TSFM model used to compare may use a different number of profiles. In fact, Figure 10 suggests that $K=20$ is the most suitable value in the VQ-TSFM context. On the other hand, the Informer approach does not have this parameter because its latent representation are real-valued embeddings.
  • Number of samples in multinomial distribution, $S$. In the multinomial approach, $S$ represents the number of samples that are drawn from the posterior distribution of the latent variables at each time step. A larger value will adapt better to the latent profiles but also complicates the detection task of the CPD. As evidenced by the results shown in 9(b), $S=5$ provides the most suitable sample size for our task.
@@
  • Size of the temporal window, $W$. The CPD model focuses on a temporal frame to assess whether its predictions are successful or not. For example, for each true event, a true positive is returned if an alarm was given by the model within the temporal window previous to that event. If the CPD did not predict any change, then a false negative is counted. This window parameter allows therefore to select how long in advance we aim to predict suicide events. We chose a prediction period of one week ($W=7$ days), which obtained a high AUC in our analysis (see 9(c)) and is brief enough to serve as short-term prediction.
@@
    • MAP ratio (default) → based on the MAP estimates of the run length, an alarm is returned if the ratio of current $r_{t}$ over the previous day $r_{t-1}$ is below the threshold:
      $\frac{r_{t}}{r_{t-1}}<\tau$
    • MAP difference → based on the MAP estimates of the run length, an alarm is returned if the difference between current $r_{t}$ and previous $r_{t-1}$ is above the threshold:
      $r_{t}-r_{t-1}>\tau$
    • Cumulative sum → based on the cumulative probability of the run length of previous days (within the specified window of size $W$), an alarm is returned if this sum is above the threshold:
      $\sum\limits_{i=0}^{W}r_{t-i}>\tau$